### v1

@@ 5.1.4 Open questions
Several questions remain still open about the (un-)decidability of (quantum) many-body properties. An important one is what remains undecidable (if anything) at finite temperature, or under other sources of noise such as disorder in the interactions (see Section 6.3).
@@ 5.2 Quantum Information
In Wolf et al. (2011) it is argued that, due to the use of quantifiers over the integers in the first-order logic formulation of many interesting quantities in quantum information theory, it could happen that they are uncomputable in general. By now, this is known to be true in the following cases:
@@ 5.2.1 Measurements and channels
The first contribution in this direction, due to Van den Nest and Briegel (2008), is a connection between measurement-based quantum computation and undecidability.
In the paradigm of measurement-based quantum computation designed by Raussendorf and Briegel (2001), one starts with a fixed universal resource quantum state. Then the desired computation is implemented by measuring sequentially each qubit of the resource state in a basis that depends on the quantum circuit one wants to implement.
It is known that there exist resource states for which measurement-based quantum computation is as powerful as the circuit model of computation. A central question in this area is then to understand which features of the resource state yield universal quantum computational power (see Gross et al. (2007) for several results in this direction).
@@
The question analyzed in Van den Nest and Briegel (2008) is the following: which property of a graph makes measurement-based computation on the associated graph state classically simulable. Surprisingly there is a clean answer: precisely those graphs whose associated logic, in a precise sense defined by Courcelle and Oum (2007), is decidable (proved through a reduction to tiling problems Seese (1991)).
A further connection between measurements and undecidability was made by Brukner (2009); Paterek et al. (2010) where a link was established between randomness of outcomes of certain quantum measurements and logical independence of propositions from a given set of axioms. A somewhat related direction connecting randomness and computability culminated in the recent result by Trejo et al. (2023) (see references therein for previous work). A quantum random number generator is proposed for which, under reasonable hypotheses, no algorithm computing any output bit can be proved correct. Experimental support is also provided.
Another connection between undecidability and measurement can be traced back to Eisert et al. (2012), where an interpretation of matrix mortality (see section 3.4.3) is given in terms of quantum measurements. The idea is the following. Any finite number of matrices $A_{1},\ldots A_{m}$ represents a quantum measurement if we interpret them as Kraus operators of a quantum channel. That is, a measurement on state $\rho$ has outcome $j\in\{1,\ldots,m\}$ with probability ${\rm tr}(A_{j}\rho A_{j}^{\dagger})$ and, in such a case, the final state becomes
@@
If we repeat the same measurement sequentially, is there a choice of outcomes that will never occur? This is the quantum measurement occurrence problem defined by Eisert et al. (2012). Note that this is equivalent to asking whether there exist $N\in\mathbb{N}$ and $j_{1},\ldots j_{N}\in\{1,\ldots,m\}$ so that
@@
As observed in Wolf et al. (2011), using the work of Hirvensalo (2007), matrix mortality problems are also undecidable if one restricts to quantum channels [^23] Indeed classical channels suffices.. The associated result is the reachability problem, which is a problem in quantum control theory. Given an initial state $\rho$, a target state $|\psi\rangle$, a parameter $10<\lambda<$ 1, and a set of quantum controls $T_{1},\cdots T_{n}$ (that are quantum channels), it is undecidable to know whether there is a choice of controls so that
@@
This connection between quantum control and undecidability was explored further in Bondar and Pechen (2020).
Gimbert and Oualhadj (2010) gave a result about the undecidability of control problems that is much stronger than (3). By a gap amplification technique, they show that, even with the promise that
@@
    Either $\langle\psi|T_{i_{N}}\cdot T_{i_{1}}(\rho)|\psi\rangle<\frac{1}{2}$ for all $N$ and all $i_{1},\cdots i_{N}$;
@@
This result was used by Elkouss and Pérez-García (2018) to show that given a classical channel with memory (the result trivially holds then for the quantum case) and the promise that its capacity is $1$ or $<\frac{1}{2}$, it is undecidable to know which is the case. This was generalized later by Boche et al. (2020), with completely different techniques, to cover the minimal possible channel size: one bit of input, one bit of output, one bit of memory.
@@
Boche and Mönich (2021) prove that the bandwidth of a bandlimited signal is uncomputable. The relevance of this result stems from the fact that, according to the Shannon sampling theorem, the bandwidth determines the minimum sampling rate which is required for a perfect reconstruction.
Boche and Deppe (2021) prove that the zero-error capacity of a memoryless channel is uncomputable. To define the zero-error capacity, there is no probability of error allowed in the transmission of information, as opposed to the standard setup where an error is tolerated, as long as it vanishes asymptotically (in the number channel uses).
However, this result comes with a caveat that is worth mentioning as it is not uncommon to results about “uncomputable” functions on spaces over the field of computable real numbers. It is a jarring fact about the field of computable numbers that it is not computably ordered (Rice (1954)). That is, statements like $x=0$ or $x>1$ are undecidable for computable real numbers in general. As a consequence, every discontinuous function, even a simple step function, becomes uncomputable. In fact, this is precisely the case for the zero-error capacity—it is discontinuous by the nature of its definition.
In such a case one could argue that the uncomputability might be more a statement about how the input is presented than a signature of any quality of the considered property that goes beyond discontinuity.
To circumvent this issue, most of the results mentioned in this review work with computably ordered subfields like the fields of rational or algebraic numbers.
@@ 5.2.2 Quantum non-locality
Bell inequalities date back to the famous work of Einstein et al. (1935), where Einstein, Podolsky and Rosen raise an apparent dilemma about the completeness of quantum mechanics as a model for Nature. 30 years later, Bell (1964) proposed an idealized experiment to test that quantum mechanics is in fact not compatible with a classical explanation of Nature in terms of local hidden variables. A series of experiments initiated by J. Clauser, A. Aspect and co-authors in the 70’s culminated in the 1982 experiment led by A. Aspect A. Aspect (1981); Aspect et al. (1982) which removed the locality loophole, and is widely wiewed as the first to convincingly demonstrating Bell inequality violation and separate quantum physics from any local hidden variable theory. In 2015 an arguably fully ‘loophole-free’ experiment that also removed certain other loopholes was performed in Delft by Hensen et al. (2015), and independently also in Vienna Giustina et al. (2015) and NIST Shalm et al. (2015), again obtaining the predicted violations of Bell inequalities. At the practical level, violations of Bell inequalities are behind some important applications of quantum information, such as quantum cryptography and the secure generation of random numbers.
@@
The game $G$ is then given by the pair $(\pi,V)$ and its size is $\max\{N,M\}$. The value of the game is the optimal probability of winning, optimized among all allowed strategies [^24] There are different possible variants. For instance, one can also take $V(a,b,x,y)$ real-valued (interpret it e.g. as the amount of money the players win if they answer $(a,b)$ on input $(x,y)$) and the value of the game would then be the maximum expected gain, when optimized over all possible strategies; one can add more players, or more rounds of communication between referee and players; one can make questions and/or answers quantum; one can allow for some type of communication between the players once the game starts; …. We refer to Buhrman et al. (2010); Palazuelos and Vidick (2016) for more details on the topic..
@@
$p(ab|xy)=\sum_{\lambda}p_{\lambda}p^{A}(a|x\lambda)p^{B}(y|\lambda)$ (4)
@@
The classical value of the game $\omega(G)$ is the optimal probability of wining if the players use classical strategies. That is
@@
There is however an important subtlety here regarding how the spatial separation between Alice and Bob is modelled. In principle, this gives rise to two different types of strategies.
@@
where $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ are finite dimensional Hilbert spaces for Alice and Bob, respectively, $\rho_{AB}$ is a mixed state on $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, and $E^{x}_{a}$ (resp. $F^{y}_{b}$) is a generalized measurement for Alice (resp. Bob). That is, $E^{x}_{a}$ are positive semidefinite operators acting on $\mathcal{H}_{A}$ so that $\sum_{a}E^{x}_{a}={\rm Id}$ for all $x$.
The entangled value $\omega^{*}(G)$ refers to those types of strategies.
@@
Bell inequalities deal with the first of those inequalities $\omega(G)\leq\omega^{*}(G)$ A Bell inequality violation in the language of non-local games is nothing but a game $G$ for which the inequality is strict $\omega(G)<\omega^{*}(G)$, meaning that quantum strategies are strictly better than classical ones for that particular game.
@@
In a breakthrough paper Slofstra (2020) showed that it is undecidable to know if $\omega^{*}_{c}(G)=1$ for a given game $G$. In his subsequent paper Slofstra (2019) showed the same result for $\omega^{*}(G)$. The proof goes via a clever mapping between finitely presented groups and a class of non-local games. The results are then obtained by invoking the undecidability of the word problem for groups (see section 3.4.4). Slofstra’s results were used later by Fritz (2021) to show that quantum logic is undecidable. They were also used by Mančinska and Roberson (2020) to give a striking undecidability result in the context of graph theory. Mančinska and Roberson show that, given two graphs $G,H$, knowing whether there exists a planar graph $K$ for which ${\rm hom}(K,G)\not={\rm hom}(K,H)$ is an undecidable problem. Here ${\rm hom}(K,G)$ is the number of graph homomorphisms from $K$ to $G$. The reason underlying this undecidability is indeed the main result of Mančinska and Roberson (2020): ${\rm hom}(K,G)={\rm hom}(K,H)$ for all planar graphs $K$ if and only if $G$ and $H$ are quantum isomorphic (see Atserias et al. (2019) for the formal definition). This in turn can be characterized by the existence of a perfect commuting strategy for a particular non-local game defined in terms of $G$ and $H$. The result then can be obtained as a reduction from Slofstra (2020). Fu et al. (2021) strengthen the results of Slofstra (2019) and Slofstra (2020) by showing that undecidability is kept even for a particular fixed game size.
@@
Finally, in a breakthrough paper, Ji et al. (2020) managed to prove that, even under the promise that $\omega^{*}(G)=1$ or $\omega^{*}(G)\leq 1/2$, it is undecidable to know which is the case. This has the remarkable consequence that there exist games for which $\omega^{*}(G)\not=\omega_{c}^{*}(G)$, giving a negative answer to Tsirelson’s problem. The way to obtain this consequence is by using undecidability as a proof ingredient, as already commented in section 5.1.3. One considers an algorithm that interleaves upper bounds to $\omega^{*}_{c}(G)$ by using the semidefinite-program-hierachy of Navascués et al. (2008), with lower bounds to $\omega^{*}(G)$ by brute-force computing the supremum in (6) for larger and larger Hilbert space dimensions. If $\omega^{*}(G)=\omega^{*}_{c}(G)$, on input a game for which $\omega^{*}(G)=1$ or $\omega^{*}(G)\leq 1/2$, such an algorithm would always halt and give the right answer, making the problem decidable, which is the desired contradiction. We refer to section 8 in Fitzsimons et al. (2019) for a detailed explanation.
@@
Using a connection made by Junge et al. (2011) and Fritz (2012) between Tsirelson’s problem and Connes’ embedding problem, the result of Ji et al. (2020) implies also a negative solution to Connes’ embedding problem. The latter was major open problem in the theories of von Nuemann algebras and $C^{*}$-algebras that is finally solved using this completely unexpected approach based on undecidability.
@@
Another consequence of Ji et al. (2020) is the undecidability of the membership problem for the set of quantum contextual behaviors. This is obtained in Wright and Farkas (2023) via the construction of an invertible map between non-locality and contextuality scenarios.
@@
In a slightly different direction, Bendersky et al. (2017) analyze the maximum set of strategies $p(ab|xy)$ that is compatible with special relativity, in the sense that the output $a$ (resp. $b$) cannot contain information about the input $y$ (resp. $x$). That is, $\sum_{b}p(ab|xy)$ is independent of $y$ and $\sum_{a}p(ab|xy)$ is independent of $x$. The set of such strategies is called non-signaling and it contains all quantum ones.
@@
By using undecidability results about learnability of functions (in turn reduced to the halting problem by Zeugmann and Zilles (2008)), Bendersky et al. (2017) prove the that non-signaling deterministic models for strategies $p(ab|xy)$ have to be uncomputable.

### v2

@@ 5.1.4 Open questions
Several questions remain still open about the (un-)decidability of (quantum) many-body properties. An important one is what remains undecidable (if anything) at finite temperature, or under other sources of noise such as disorder in the interactions (see Section 6.3). For instance, to the best of our knowledge the following question is still open: given a translational invariant Hamiltonian, is it decidable or undecidable to determine if the correlations in the associated thermal state decay exponentially with the distance?
@@ 5.2 Quantum Information
It is argued by Wolf et al. (2011) [140] that, due to the use of quantifiers over the integers in the first-order logic formulation of many interesting quantities in quantum information theory, it could happen that they are uncomputable in general. By now, this is known to be true in the following cases:
@@ 5.2.1 Measurements and channels
The first contribution in this direction, due to Van den Nest and Briegel (2008) [141], is a connection between measurement-based quantum computation and undecidability.
In the paradigm of measurement-based quantum computation designed by Raussendorf and Briegel [142], one starts with a fixed universal resource quantum state. Then the desired computation is implemented by measuring sequentially each qubit of the resource state in a basis that depends on the quantum circuit one wants to implement.
It is known that there exist resource states for which measurement-based quantum computation is as powerful as the circuit model of computation. A central question in this area is then to understand which features of the resource state yield universal quantum computational power (see [143] for several results in this direction).
@@
The question analyzed in [141] is the following: which property of a graph makes measurement-based computation on the associated graph state classically simulable. Surprisingly there is a clean answer: precisely those graphs whose associated logic, in a precise sense defined in [144], is decidable (proved through a reduction to tiling problems [145]).
A further connection between measurements and undecidability was made by Brukner (2009) and Paterek et al. (2010) [42, 43] where a link was established between randomness of outcomes of certain quantum measurements and logical independence of propositions from a given set of axioms. A somewhat related direction connecting randomness and computability culminated in the recent result by Trejo et al. (2023) [44] (see references therein for previous work). A quantum random number generator is proposed for which, under reasonable hypotheses, no algorithm computing any output bit can be proved correct. Experimental support is also provided.
Another connection between undecidability and quantum measurements can be traced back to the work of Eisert et al. (2012) [70], where an interpretation of matrix mortality (see section 3.4.3) is given in terms of quantum measurements. The idea is the following. Any finite number of matrices $A_{1},\ldots A_{m}$ represents a quantum measurement if we interpret them as Kraus operators of a quantum channel. That is, a measurement on state $\rho$ has outcome $j\in\{1,\ldots,m\}$ with probability ${\rm tr}(A_{j}\rho A_{j}^{\dagger})$ and, in such a case, the final state becomes
@@
If we repeat the same measurement sequentially, is there a choice of outcomes that will never occur? This is the quantum measurement occurrence problem defined in [70]. Note that this is equivalent to asking whether there exist $N\in\mathbb{N}$ and $j_{1},\ldots j_{N}\in\{1,\ldots,m\}$ so that
@@
As observed by Wolf et al. (2011) [140], using the work of Hirvensalo (2007) [146], matrix mortality problems are also undecidable if one restricts to quantum channels [^24] Indeed classical channels suffices.. The associated result is the reachability problem, which is a problem in quantum control theory. Given an initial state $\rho$, a target state $|\psi\rangle$, a parameter $10<\lambda<$ 1, and a set of quantum controls $T_{1},\cdots T_{n}$ (that are quantum channels), it is undecidable to know whether there is a choice of controls so that
@@
This connection between quantum control and undecidability was explored further by Bondar and Pechen (2020) [63].
Gimbert and Oualhadj (2010) [147] gave a result about the undecidability of control problems that is much stronger than (4). By a gap amplification technique, they show that, even with the promise that
@@
    either $\langle\psi|T_{i_{N}}\cdots T_{i_{1}}(\rho)|\psi\rangle<\frac{1}{2}$ for all $N$ and all $i_{1},\cdots i_{N}$;
@@
This result was used by Elkouss and Pérez-García (2018) [148] to show that given a classical channel with memory (the result trivially holds then for the quantum case) and the promise that its capacity is $1$ or $<\frac{1}{2}$, it is undecidable to know which is the case. This was generalized later by Boche et al. (2020) [149], with completely different techniques, to cover the minimal possible channel size: one bit of input, one bit of output, one bit of memory.
@@
Boche and Mönich (2021) [150] proved that the bandwidth of a bandlimited signal is uncomputable. The relevance of this result stems from the fact that, according to the Shannon sampling theorem, the bandwidth determines the minimum sampling rate which is required for a perfect reconstruction.
Boche and Deppe (2021) [151] proved that the zero-error capacity of a memoryless channel is uncomputable. To define the zero-error capacity, there is no probability of error allowed in the transmission of information, as opposed to the standard setup where an error is tolerated, as long as it vanishes asymptotically (in the number channel uses).
However, this result comes with a caveat that is worth mentioning as it is not uncommon to results about “uncomputable” functions on spaces over the field of computable real numbers. It is a jarring fact about the field of computable numbers that it is not computably ordered [152]. That is, statements like $x=0$ or $x>1$ are undecidable for computable real numbers in general. As a consequence, every discontinuous function, even a simple step function, becomes uncomputable. In fact, this is precisely the case for the zero-error capacity—it is discontinuous by the nature of its definition.
In such a case, one could argue that the uncomputability might be more a statement about how the input is presented than a signature of any quality of the considered property that goes beyond discontinuity.
To circumvent this issue, most of the results mentioned in this review work with computably ordered subfields like the fields of rational or algebraic numbers. In the case of the zero-error capacity, one could modify the question and, for instance, consider the decision problem of whether the capacity is larger than a given value in cases where everything takes place within a computably ordered field (e.g. the algebraic numbers). To the best of our knowledge, it is currently open whether this problem is (un-)decidable.
@@ 5.2.2 Quantum non-locality
Bell inequalities date back to the famous work of [153], where Einstein, Podolsky and Rosen raise an apparent dilemma about the completeness of quantum mechanics as a model for nature. 30 years later, Bell [154] proposed an idealized experiment to test that quantum mechanics is in fact not compatible with a classical explanation of Nature in terms of local hidden variables. Implementing this experiment was a challenging task since one needed to avoid some crucial mismatches –called loopholes– between the required ideal setup and the actual experiment. The two main loopholes to deal with were the locality loophole and the detection loophole  [^25] There are many other loopholes one could consider, as one increases the level of scrutiny regarding potential ways nature might conspire to invalidate the Bell experiment. As A.  Aspect says [155]: “Yet no experiment, as ideal as it is, can be said to be totally loophole-free […]. Since all events have a common past if we go back far enough in time –possibly to the Big Bang– any observed correlation could be justified by invoking such an explanation.”.. The locality loophole refers to the need to avoid causality relations between the spatially separated measurement setups of the experiment. Solving the locality loophole requires fast random measurement choices in order to avoid any causal influence of the measurement choice at one side on the outcome at the other side. The detection loophole refers to the problem that only a small fraction of the photons are actually detected and that this fraction could depend on the configuration of the measurement setups.
A series of experiments initiated by J. Clauser, A. Aspect and co-authors in the 70’s culminated in the 1982 experiment led by A. Aspect [156, 157] which removed the locality loophole, and is widely viewed as the first to convincingly demonstrating Bell inequality violation and separate quantum physics from any local hidden variable theory. Finally, in 2015 an experiment removing simultaneously the locality and detection loopholes was performed in Delft [158], and independently also in Vienna [159] and NIST [160], again obtaining the predicted violations of Bell inequalities.
At the practical level, violations of Bell inequalities are behind some important applications of quantum information, such as quantum cryptography [161] and the secure generation of random numbers [162, 163].
@@
The game $G$ is then given by the pair $(\pi,V)$ and its size is $\max\{N,M\}$. The value of the game is the optimal probability of winning, optimized among all allowed strategies [^26] There are different possible variants. For instance, one can also take $V(a,b,x,y)$ to be real-valued (interpret it e.g. as the amount of money the players win if they answer $(a,b)$ on input $(x,y)$) and the value of the game would then be the maximum expected gain, when optimized over all possible strategies; one can add more players, or more rounds of communication between referee and players; one can make questions and/or answers quantum; one can allow for some type of communication between the players once the game starts; …. We refer to [164, 165] for more details on the topic..
@@
$p(ab|xy)=\sum_{\lambda}p_{\lambda}p^{A}(a|x\lambda)p^{B}(b|y\lambda)$ (5)
@@
The classical value of the game $\omega(G)$ is the optimal probability of winning if the players use classical strategies. That is
@@
There is however an important subtlety here regarding how the spatial separation between Alice and Bob is modeled. In principle, this gives rise to two different types of strategies.
@@
where $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ are finite dimensional Hilbert spaces for Alice and Bob, respectively, $\rho_{AB}$ is a mixed state on $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, and $E^{x}_{a}$ (resp. $F^{y}_{b}$) is a generalized measurement for Alice (resp. Bob). That is, $E^{x}_{a}$ are positive semidefinite operators acting on $\mathcal{H}_{A}$ so that $\sum_{a}E^{x}_{a}={\rm Id}$ for all $x$. The entangled value $\omega^{*}(G)$ refers to those types of strategies and is defined as follows:
@@
Bell inequalities deal with the first of those inequalities $\omega(G)\leq\omega^{*}(G)$. A Bell inequality violation in the language of non-local games is nothing but a game $G$ for which the inequality is strict $\omega(G)<\omega^{*}(G)$, meaning that quantum strategies are strictly better than classical ones for that particular game.
@@
In a breakthrough paper Slofstra (2020) [75] showed that it is undecidable to know if $\omega^{*}_{c}(G)=1$ for a given game $G$. In his subsequent paper Slofstra (2019) [74] showed the same result for $\omega^{*}(G)$. The proof goes via a clever mapping between finitely presented groups and a class of non-local games. The results are then obtained by invoking the undecidability of the word problem for groups (see section 3.4.4). Slofstra’s results were used later by Fritz (2021) [76] to show that quantum logic is undecidable. They were also used by Mančinska and Roberson (2020) [77] to give a striking undecidability result in the context of graph theory. They show that, given two graphs $G,H$, knowing whether there exists a planar graph $K$ for which ${\rm hom}(K,G)\not={\rm hom}(K,H)$ is an undecidable problem. Here ${\rm hom}(K,G)$ is the number of graph homomorphisms from $K$ to $G$. The reason underlying this undecidability is indeed the main result of [77]: ${\rm hom}(K,G)={\rm hom}(K,H)$ for all planar graphs $K$ if and only if $G$ and $H$ are quantum isomorphic (see [166] for the formal definition). This in turn can be characterized by the existence of a perfect commuting strategy for a particular non-local game defined in terms of $G$ and $H$. The result then can be obtained as a reduction from [75]. Fu et al. (2021) [78] strengthened the results of [74, 75] by showing that undecidability is kept even for a particular fixed game size.
@@
Finally, in a breakthrough paper, Ji et al. (2020) [167] managed to prove that, even under the promise that $\omega^{*}(G)=1$ or $\omega^{*}(G)\leq 1/2$, it is undecidable to know which is the case. This has the remarkable consequence that there exist games for which $\omega^{*}(G)\not=\omega_{c}^{*}(G)$, giving a negative answer to Tsirelson’s problem. The way to obtain this consequence is by using undecidability as a proof ingredient, as already commented in section 5.1.3. One considers an algorithm that interleaves upper bounds to $\omega^{*}_{c}(G)$ by using the semidefinite-program-hierarchy of Navascués et al. (2008) [168], with lower bounds to $\omega^{*}(G)$ by brute-force computing the supremum in (7) for larger and larger Hilbert space dimensions. If $\omega^{*}(G)=\omega^{*}_{c}(G)$, on input a game for which $\omega^{*}(G)=1$ or $\omega^{*}(G)\leq 1/2$, such an algorithm would always halt and give the right answer, making the problem decidable, which is the desired contradiction. We refer to section 8 in [169] for a detailed explanation.
@@
Using a connection made in [170] and [171] between Tsirelson’s problem and Connes’ embedding problem, the result in [167] implies also a negative solution to Connes’ embedding problem. The latter was a major open problem in the theories of von Nuemann algebras and $C^{*}$-algebras that is finally solved using this completely unexpected approach based on undecidability.
@@
Another consequence of Ji et al. (2020) [167] is the undecidability of the membership problem for the set of quantum contextual behaviors. This was obtained in Wright and Farkas (2023) [172] via the construction of an invertible map between non-locality and contextuality scenarios.
@@
In a slightly different direction, Bendersky et al. (2017) [173] analyzed the maximum set of strategies $p(ab|xy)$ that is compatible with special relativity, in the sense that the output $a$ (resp. $b$) cannot contain information about the input $y$ (resp. $x$). That is, $\sum_{b}p(ab|xy)$ is independent of $y$ and $\sum_{a}p(ab|xy)$ is independent of $x$. The set of such strategies is called non-signaling and it contains all quantum ones.
@@
By using undecidability results about learnability of functions (in turn reduced to the halting problem in [174]), it is proven in [173] that non-signaling deterministic models for strategies $p(ab|xy)$ have to be uncomputable.